A steel ballA of mass 20 kg moving with a speed of 2m/s collide with another ball B of mass 10 kg which is initially at rest. After collision, a move with a speed of 1m/s at an angle of 30° with its original direction of motion. Determine the final velocity
1
Expert's answer
2020-03-11T10:28:34-0400

During collision total momentum of the system remains conserved

"P_i=P_f"


in x direction

"20\times2 +0=20\times1cos 30^o+10\times v_{x}\\v_x=(40-10\sqrt3)10=(4-\sqrt3)m/s"


in y direction

"0=20\times1sin30^o-10v_y\\v_y=1 m/s"